# Adaptive Alpha Weighting with PPO: Enhancing Prompt-Based LLM-Generated Alphas in Quant Trading

### v1

@@ 3.2.1 Prompt-Based LLM for Alpha Generation
The model used to generate formulaic alphas is a prompt-based LLM, specifically deepseek-r1-distill-llama-70b. This LLM is from the American AI company Groq [^2] https://groq.com/. This version of DeepSeek model has been shown to generate competitive results without fine-tuning compared to other LLMs such as LLaMA and Gemma in previous study Chen2025 . This model (deepseek-r1-distill-llama-70b) uses knowledge distillation, a process in which a smaller model (the student) learns from the output of a larger, more complex model (the teacher). In this case, the LLaMA-70B model serves as the teacher, and the student model is trained to replicate its behavior. The teacher model generates predictions or soft targets, which guide the student in learning more efficiently. Knowledge distillation allows the student model to retain much of the performance of the teacher model while being more lightweight, making it suitable for deployment in environments with limited resources hinton2015distillingknowledgeneuralnetwork .
@@
For example, one formula might combine momentum indicators with moving averages, while another might include sentiment scores to account for market news. Fifty formulas are generated for the forecasting of stock returns of all companies.
@@
You are a quantitative trader. Generate 50 alpha formulas using the given stock features:
{features}.
@@
Return only the formulas in Python syntax, using variables like C_t (Close), O_t (Open),
V_t (Volume), S_t (Sentiment), and standard indicators (SMA, Momentum).
@@ 3.2.2 PPO for Alpha Weight Optimization
Table 3: Alpha Trading Algorithm
Algorithm: Alpha Trading with Adaptive Risk Controls
@@
 Price data $\mathcal{D}=\{OHLCV_{t},\text{FutureReturn}_{t}\}_{t=1}^{T}$
@@
 Transaction cost $\lambda=0.001$
@@
 Performance metrics $\mathcal{M}$
@@
1. Preprocess data:
$\alpha_{i}\leftarrow\text{StandardScaler.fit\_transform}(\alpha_{i}),\ \forall i\in 1:50$
$\text{MA}_{20}\leftarrow\text{rolling\_mean}(\text{Close},20)$
$\text{MA}_{100}\leftarrow\text{rolling\_mean}(\text{Close},100)$
$\text{regime}\leftarrow\mathbb{I}(\text{MA}_{20}>\text{MA}_{100})$
$\sigma_{\text{daily}}\leftarrow\text{rolling\_std}(\text{FutureReturn},63)$
$\sigma_{\text{annual}}\leftarrow\sigma_{\text{daily}}\times\sqrt{252}$
$\tau^{\text{upper}}\leftarrow\text{rolling\_quantile}(\text{Close},0.75,126)$
$\tau^{\text{lower}}\leftarrow\text{rolling\_quantile}(\text{Close},0.25,126)$
2. Initialize tracking: $V\leftarrow 1.0,\ p\leftarrow 0,\ \text{peak}\leftarrow 1.0$
@@ 1. Preprocess data:
 4. Agent Decision:
  $\mathbf{w}_{t}\leftarrow\text{model.predict}(s_{t})$
@@ 4. Agent Decision:
  $\mathbf{w}_{t}^{\text{norm}}\leftarrow\frac{\tilde{\mathbf{w}}_{t}}{\|\tilde{\mathbf{w}}_{t}\|_{1}+\epsilon},\quad\epsilon=10^{-8}$
 5. Composite Signal:
  $\alpha^{\text{composite}}_{t}\leftarrow\sum_{i=1}^{50}w_{t}[i]\cdot\alpha_{i,t}$
 6. Position Sizing:
@@
[…]
  Vol scaling: $v_{t}\leftarrow\min(2,\sigma_{\text{target}}/\sigma^{\text{annual}}_{t})$
@@
 7. Execute Trade:
  $\text{cost}\leftarrow\lambda\cdot|p_{t}-p_{t-1}|$
  $r_{t}\leftarrow p_{t}\cdot\text{FutureReturn}_{t}-\text{cost}$
  $V_{t}\leftarrow V_{t-1}\cdot(1+r_{t})$
@@ 8. Update State:
Performance Evaluation:
9. Compute metrics:
$\text{Sharpe}\leftarrow\frac{\mathbb{E}[R]}{\sigma_{R}}\times\sqrt{252}$
$\text{MaxDD}\leftarrow\min(\text{Cumulative}/\text{cumulative\_max}(\text{Cumulative})-1)$
@@ 9. Compute metrics:
$\lambda$ Transaction cost per trade 0.001
@@ 3.3 Evaluation
To evaluate the effectiveness of the proposed framework, we compare its trading performance with an equal-weighted strategy and major market indices. In the equal-weighted strategy, each LLM-generated alpha is assigned the same weight within the portfolio. The framework is considered effective if it consistently outperforms both the equal-weighted strategy and the market indices. To assess performance comprehensively, we use a range of metrics, including Information Coefficient (IC), Mutual Information, LightGBM Feature Importance, Cumulative Return, Maximum Drawdown, and Sharpe Ratio.
First, the Information Coefficient (IC) (Equation 9) is calculated for each alpha signal across the five selected companies: Apple, HSBC, Pepsi, Tencent, and Toyota. The IC measures the correlation between the predicted signals and the actual future returns. A positive IC value close to or above 0.05 indicates that the alpha has significant predictive power. IC values below 0.05 suggest weak predictive power, while negative IC values imply an inverse relationship, which may still be useful to construct contrarian strategies.
A contrarian strategy is an investment approach that opposes prevailing market trends. Contrarian investors buy when others are selling and sell when others are buying, anticipating price reversals dreman1998contrarian . In this context, a negative IC indicates that the signal moves in the opposite direction of return. Although the signal does not directly predict returns, it can still be valuable for contrarian strategies, where trading against the signal’s direction is intentional.
@@
The Cumulative Return of each company’s portfolio, optimized by the Proximal Policy Optimization (PPO) algorithm, is compared to its respective market benchmark (e.g., S&P 500 for Apple and Pepsi), Hang Seng Index for HSBC and Tencent, and Nikkei 225 for Toyota. This comparison assesses whether the RL-optimized alpha weights can outperform traditional market indices over the evaluation period.
@@ 4 Results
The following formulaic alphas (Table 5) were generated by the LLM. These alphas are grouped by their main focus, including momentum, sentiment, volume, market indices, technical indicators, moving averages, volatility, and various combinations. These alphas incorporate price, volume, and sentiment features to capture diverse trading signals.
@@ 4.1 MI and IC Analysis
Figure 4 presents the mutual information (MI) and the information coefficient (IC) of the fifty LLM-generated alphas in five companies. In the figure, the blue bars indicate the MI values, while the red bars represent the IC values. Together, they offer insight into both the statistical dependence between the alphas and returns (MI), and the directional predictive power (IC).
@@ 4.2 Feature Importance
Figure 5 displays the LightGBM feature importance scores for each LLM-generated alpha across the five selected companies. The results indicate that Alpha 25 stands out as the most influential feature for Apple, Toyota, and Pepsi, suggesting that this alpha consistently contributes valuable information to the prediction models for these companies. In contrast, Alpha 21 holds the highest importance for HSBC and Tencent, implying that different alphas capture distinct patterns depending on the company. On the other hand, Alpha 39 appears to have minimal importance across all five companies, indicating limited predictive relevance in most cases. While there are differences in which alphas are prioritized, the overall distribution of importance scores does not show extreme variation, suggesting a relatively balanced reliance on multiple alphas rather than dominance by a few. This diversity highlights that the predictive power of each alpha may depend heavily on the underlying characteristics of each stock.
4.3 Comparison of PPO-Adjusted Strategy and Market Benchmark

### v2

@@ 3.2.1 Prompt-Based LLM for Alpha Generation
The model used to generate formulaic alphas is a prompt-based LLM, specifically deepseek-r1-distill-llama-70b. This LLM is provided by American AI company Groq [^2] https://groq.com/ for deployment and inference. This version of DeepSeek model has been shown to generate competitive results without fine-tuning compared to other LLMs such as LLaMA and Gemma in previous study Chen (2025e). This model (deepseek-r1-distill-llama-70b) uses knowledge distillation, a process in which a smaller model (the student) learns from the output of a larger, more complex model (the teacher). In this case, the LLaMA-70B model serves as the teacher, and the student model is trained to replicate its behavior. The teacher model generates predictions or soft targets, which guide the student in learning more efficiently. Knowledge distillation allows the student model to retain much of the performance of the teacher model while being more lightweight, making it suitable for deployment in environments with limited resources Hinton et al. (2015).
@@
For example, one formula might combine momentum indicators with moving averages, while another might include sentiment scores to account for market news. Fifty formulas are generated for the forecasting of stock returns of all companies. To avoid look-ahead bias, only training data is used to generate formulaic alphas.
@@
You are a quantitative trader. Generate 50 alpha formulas using the given stock features: {features}.
@@
Return only the formulas in Python syntax, using variables like C_t (Close), O_t (Open), V_t (Volume), S_t (Sentiment), and standard indicators (SMA, Momentum).
@@ 3.2.2 PPO for Alpha Weight Optimization
We further elaborate on the definition and role of the market regime in our framework. To incorporate regime-aware risk control, we introduce a soft penalty when the trading position is inconsistent with the prevailing market regime. The market regime indicator $\text{regime}_{t}\in\{0,1\}$ is defined based on the moving-average crossover, where $\text{regime}_{t}=1$ denotes a bullish regime (20-day MA above 100-day MA) and $\text{regime}_{t}=0$ indicates a bearish regime.
A regime violation occurs when the sign of the position conflicts with the regime signal, i.e.,
$\mathbb{I}_{t}^{\text{regime}}=\begin{cases}1,&(\text{regime}_{t}=0\ \wedge\ p_{t}>0)\ \text{or}\ (\text{regime}_{t}=1\ \wedge\ p_{t}<0),\\ 0,&\text{otherwise},\end{cases}$ (8)
and the corresponding regime penalty is defined as
$\mathcal{P}_{t}^{\text{regime}}=\lambda_{\text{reg}}\cdot|p_{t}|\cdot\mathbb{I}_{t}^{\text{regime}},$ (9)
where $\lambda_{\text{reg}}$ controls the strength of the penalty. In this study, $\lambda_{\text{reg}}$ equals to 0.05.
@@
Table 3: Alpha Trading Algorithm with Adaptive Risk Controls
Algorithm: PPO-based Alpha Trading Environment
@@
 Price data $\mathcal{D}=\{OHLCV_{t},R^{\text{future}}_{t}\}_{t=1}^{T}$
@@
 Transaction cost coefficient $\lambda$
 Regime penalty coefficient $\lambda_{\text{reg}}$
@@
 Portfolio value series $\{V_{t}\}_{t=1}^{T}$ and performance metrics $\mathcal{M}$
@@
1. Feature preprocessing:
$\alpha_{i,t}\leftarrow\text{StandardScaler}(\alpha_{i,t}),\ \forall i\in\{1,\dots,50\}$
$\text{MA}_{20,t}\leftarrow\text{MA}(\text{Close}_{t},20),\quad\text{MA}_{100,t}\leftarrow\text{MA}(\text{Close}_{t},100)$
$\text{regime}_{t}\leftarrow\mathbb{I}(\text{MA}_{20,t}>\text{MA}_{100,t})$
$\sigma^{\text{daily}}_{t}\leftarrow\text{Std}(R^{\text{future}}_{t-62:t})$
$\sigma^{\text{annual}}_{t}\leftarrow\sigma^{\text{daily}}_{t}\sqrt{252}$
$\tau^{\text{upper}}_{t}\leftarrow Q_{0.75}(\text{Close}_{t-126:t}),\;\tau^{\text{lower}}_{t}\leftarrow Q_{0.25}(\text{Close}_{t-126:t})$
2. State initialization:
$V_{0}\leftarrow 1.0,\;p_{0}\leftarrow 0,\;\text{peak}\leftarrow 1.0$
@@ 2. State initialization:
 4. Policy action:
  $\mathbf{w}_{t}\leftarrow\pi_{\theta}(s_{t})$
@@ 4. Policy action:
  $\mathbf{w}_{t}^{\text{norm}}\leftarrow\tilde{\mathbf{w}}_{t}/(\|\tilde{\mathbf{w}}_{t}\|_{1}+\epsilon)$
 5. Composite alpha construction:
  $\alpha^{\text{composite}}_{t}\leftarrow\sum_{i=1}^{50}w_{t}^{\text{norm}}[i]\cdot\alpha_{i,t}$
 6. Signal-based position sizing:
@@
  Volatility targeting:
  $v_{t}\leftarrow\min(2,\sigma_{\text{target}}/\sigma^{\text{annual}}_{t})$
@@
  Regime-aware risk penalty:
  $\mathcal{P}_{t}^{\text{regime}}\leftarrow\lambda_{\text{reg}}|p_{t}|\cdot\mathbb{I}(\text{sign}(p_{t})\neq\text{regime}_{t})$
 7. Reward and portfolio update:
  $\text{TC}_{t}\leftarrow\lambda|p_{t}-p_{t-1}|$
  $r_{t}\leftarrow p_{t}R^{\text{future}}_{t}-\text{TC}_{t}-\mathcal{P}_{t}^{\text{regime}}$
  $V_{t}\leftarrow V_{t-1}(1+r_{t})$
@@ 8. State transition:
Evaluation:
9. Compute out-of-sample metrics (Sharpe ratio, max drawdown, turnover).
@@
$\text{TC}_{t}$ Transaction cost per trade 0.001
@@
3.3 Implementation Details
Timet1t-1ttt+1t+1 Information Set (Regime, Vol, LLM Alphas) Signal / Position $p_{t}$Holding Period ($p_{t}$)Return $r_{t+1}$realized at closeReturn Attribution: $R_{t+1}=p_{t}\cdot r_{t+1}$
At each trading day $t$, the agent makes a trading decision based solely on information available up to the end of day $t$. The information set includes regime indicators, realized volatility estimates, and standardized LLM-generated alpha signals computed from historical data only. Trades are executed at the close price of day $t$, and no information beyond day $t$ is used when determining trading positions.
The agent outputs a vector of weights over the LLM-generated alphas at day $t$, which are aggregated into a composite alpha signal. This composite signal is then mapped to a target position $p_{t}\in[-1,1]$, representing a normalized long or short exposure. Position sizing is further adjusted through a volatility-targeting mechanism to maintain a stable risk profile across different market conditions.
Trades are assumed to be executed immediately after the decision at day $t$, and the resulting position $p_{t}$ is held over the interval from day $t$ to day $t+1$. The realized strategy return is computed using the next-day return $r_{t+1}$, and the portfolio value is updated accordingly. Specifically, the strategy return at day $t+1$ is given by
$R_{t+1}=p_{t}\cdot r_{t+1},$ (11)
where transaction costs and turnover effects are modeled separately during training.
This trading protocol ensures that signals are formed strictly prior to return realization, avoiding look-ahead bias and making the strategy fully implementable in practice.
@@ 3.4 Evaluation
3.4.1 Baselines
To evaluate the effectiveness of the proposed framework, we compare its trading performance with an equal-weighted strategy, buy-and-hold strategy, random entry/exit strategy, and momentum strategy. In the equal-weighted strategy, each LLM-generated alpha is assigned the same weight within the portfolio. The following provides a brief introduction to the remaining baseline strategies.
Buy and Hold (B&H) Strategy
The Buy and Hold strategy serves as a passive benchmark representing the underlying asset’s natural market growth. Following the logic in our implementation, the strategy enters a long position at the initial timestamp of the test period and maintains this exposure until the final step. This baseline provides a reference for the "Market Beta," allowing us to determine if the RL agent’s active management adds value beyond simple market exposure.
Random Entry/Exit Strategy
To address the potential for stochastic success, we implement a Random Entry/Exit strategy that mimics the execution style of the PPO agent. This baseline is constrained by the empirical turnover and average position duration $D$ of the proposed PPO strategy. At each time step, the probability of a position switch is defined as $P(switch)=1/D$. A state transition occurs if a randomly sampled value from a uniform distribution $u\sim U(0,1)$ is less than $P$. This ensures that the random baseline maintains a turnover rate and market exposure comparable to the PPO agent. To ensure statistical reliability, we conduct 1,000 Monte Carlo simulations and report the mean and standard deviation of the resulting metrics. This allows us to verify whether the agent’s alpha is statistically distinct from a naïve, non-informative trading pattern.
Momentum (MOM) Strategy
The Momentum strategy serves as a technical baseline to evaluate whether the PPO agent can outperform a traditional trend-following heuristic. Based on the "momentum effect" observed in financial markets, this strategy calculates the return of each asset over a fixed look-back period of 10 days. The strategy enters a long position if the historical return is positive and a short position if it is negative. This baseline is essential for determining whether the RL agent is simply following prevailing price trends or if its integration of formulaic alphas enables it to identify more complex, non-linear market signals that a simple trend-following rule would miss.
3.4.2 Evaluation Metrics
To assess performance comprehensively, we use a range of metrics, including Information Coefficient (IC), Mutual Information, LightGBM Feature Importance, Cumulative Return, Maximum Drawdown, and Sharpe Ratio.
First, the Information Coefficient (IC) (Equation 12) is calculated for each alpha signal across the ten selected companies. The IC measures the correlation between the predicted signals and the actual future returns. A positive IC value close to or above 0.05 indicates that the alpha has significant predictive power. IC values below 0.05 suggest weak predictive power, while negative IC values imply an inverse relationship, which may still be useful to construct contrarian strategies.
A contrarian strategy is an investment approach that opposes prevailing market trends. Contrarian investors buy when others are selling and sell when others are buying, anticipating price reversals Dreman (1998). In this context, a negative IC indicates that the signal moves in the opposite direction of return. Although the signal does not directly predict returns, it can still be valuable for contrarian strategies, where trading against the signal’s direction is intentional.
@@ 3.4.2 Evaluation Metrics
The Cumulative Return measures the overall profitability of the trading strategy by aggregating returns over the entire evaluation period, reflecting the total gain or loss achieved through sequential trades.
@@
3.4.3 Statistical Comparison via Diebold-Mariano Test
To determine whether the performance gains of the PPO-based agent are statistically meaningful or merely a product of stochastic market noise, we utilize the Diebold-Mariano (DM) test. While traditional applications of the DM test focus on mean squared error in point forecasting, we adapt the framework to evaluate the relative profitability of trading trajectories.
We define the loss function $L(\cdot)$ as the negative daily net return of a strategy, such that $L(r_{t})=-r_{t}$. The loss differential at time $t$ is then expressed as:
$d_{t}=L(r_{t,\text{PPO}})-L(r_{t,\text{Base}})=r_{t,\text{Base}}-r_{t,\text{PPO}}$ (18)
where $r_{t,\text{PPO}}$ and $r_{t,\text{Base}}$ denote the net returns of the agent and the baseline, respectively. The null hypothesis ($H_{0}$) posits that there is no expected difference in the performance of the two strategies, effectively $\mathbb{E}[d_{t}]=0$.
The DM test statistic is calculated as:
$DM=\frac{\bar{d}}{\sqrt{\hat{\sigma}_{d}^{2}/T}}$ (19)
where $\bar{d}$ represents the sample mean of the differential series and $\hat{\sigma}_{d}^{2}$ is a consistent estimate of the variance. Given that our trading decisions occur daily with a one-step horizon, we assume a forecast lag of $h=1$. A significantly negative $DM$ statistic accompanied by a $p$-value below the threshold of $\alpha=0.05$ allows us to reject $H_{0}$, confirming that the agent’s outperformance is statistically robust.
3.4.4 Sharpe Ratio Bootstrap Test
We evaluate the risk-adjusted performance difference between the PPO strategy and the Buy-and-Hold benchmark using a bootstrap-based Sharpe ratio test. Let $\{r_{t}^{\text{PPO}}\}_{t=1}^{T}$ and $\{r_{t}^{\text{B\&H}}\}_{t=1}^{T}$ denote the daily return series of the two strategies over the test period. For the PPO strategy, daily returns are obtained by averaging across multiple independent training runs with different random seeds.
The Sharpe ratio of strategy $s\in\{\text{PPO},\text{B\&H}\}$ is defined as
$SR_{s}=\frac{\mathbb{E}\!\left[r_{t}^{s}\right]}{\sqrt{\mathrm{Var}\!\left(r_{t}^{s}\right)}}.$ (20)
We focus on the relative Sharpe ratio difference,
$\Delta SR=SR_{\text{PPO}}-SR_{\text{B\&H}}.$ (21)
To account for temporal dependence in financial returns, we adopt a block bootstrap procedure. The return series are resampled using contiguous blocks of fixed length, and for each bootstrap replication $b=1,\dots,B$, a bootstrapped Sharpe difference $\Delta SR^{(b)}$ is computed. This yields an empirical distribution $\{\Delta SR^{(b)}\}_{b=1}^{B}$.
A two-sided 95% confidence interval is constructed from the empirical quantiles of the bootstrap distribution, and the corresponding $p$-value is obtained. Statistical significance is determined by whether the confidence interval excludes zero, indicating a meaningful difference in risk-adjusted performance between the PPO strategy and Buy-and-Hold.
@@ 4 Results
4.1 LLM-Generated Formulaic Alphas
A sample of the LLM-Generated formulaic alphas is shown in Figure 5. These alphas are grouped by their main focus, including momentum, sentiment, volume, market indices, technical indicators, moving averages, volatility, and various combinations. These alphas incorporate price, volume, and sentiment features to capture diverse trading signals.
Specifically, the momentum and moving average formulas (e.g., $\alpha_{1},\alpha_{26}$) are based on the trend-following effect, where stock prices tend to keep moving in the same direction because investors often follow the crowd. Sentiment-based alphas (e.g., $\alpha_{6},\alpha_{10}$) capture how investor emotions and news can push prices away from their true value in the short term. Volume and volatility formulas (e.g., $\alpha_{11},\alpha_{36}$) use trading activity and price swings to measure market strength and risk. Additionally, the index-based and combination formulas show how local stocks are influenced by global markets, such as the S&P 500 or Nikkei 225. By turning these ideas into simple formulas, the LLM allows the PPO agent to easily find and use the connections between price, sentiment, and global economic signals.
@@ 4.2 MI and IC Analysis
Figure 3 presents the mutual information (MI) and information coefficient (IC) of fifty LLM-generated alphas for five selected companies, chosen from the full set of ten companies for illustration. In the figure, the blue bars indicate the MI values, while the red bars represent the IC values. Together, they offer insight into both the statistical dependence between the alphas and returns (MI), and the directional predictive power (IC).
@@ 4.3 Feature Importance
Figure 4 presents the LightGBM feature importance scores of each LLM-generated alpha for five selected companies, chosen from the full set of ten for illustrative purposes. The results indicate that Alpha 25 stands out as the most influential feature for Apple, Toyota, and Pepsi, suggesting that this alpha consistently contributes valuable information to the prediction models for these companies. In contrast, Alpha 21 holds the highest importance for HSBC and Tencent, implying that different alphas capture distinct patterns depending on the company. On the other hand, Alpha 39 appears to have minimal importance across all five companies, indicating limited predictive relevance in most cases. While there are differences in which alphas are prioritized, the overall distribution of importance scores does not show extreme variation, suggesting a relatively balanced reliance on multiple alphas rather than dominance by a few. This diversity highlights that the predictive power of each alpha may depend heavily on the underlying characteristics of each stock.
4.4 Comparison of PPO-Adjusted Strategy and Benchmarks
Trading performance comparison